# STRIDE: Sparse Techniques for Regression in Deep Gaussian Processes

### v1

@@ 5 Conclusion
Important future research directions are generalisations of STRIDE to non-Gaussian likelihoods, the SASPA framework might be a useful point to start [29]. Moreover, we have not analysed STRIDE in this work, techniques from [16, 1, 4, 26] might support a convergence analysis.
Limitations.
STRIDE makes deep GPR more scalable, but does not remove every limitation deep or standard GPR has. Especially the use of STRIDE with high input dimension $d$ without separate feature extraction, for instance, should still be very challenging. STRIDE profits from sparse GPR, but the more flexible deep GP models still require more computational resources.
@@
SU, ALT, and JL acknowledge the use of the HWU high-performance computing facility (DMOG) and associated support services in the completion of this work. ALT was partially supported by EPSRC grants EP/X01259X/1 and EP/Y028783/1.

### v2

@@ 5 Conclusion
Important future research directions are generalisations of STRIDE to non-Gaussian likelihoods, the SASPA framework might be a useful point to start [Qi]. Moreover, we have not analysed STRIDE in this work, techniques from [finocchio2023posterior, abraham2023deep, Burt, Osborne] might support a convergence analysis.
However, while STRIDE makes deep GPR more scalable, some limitations remain. Especially the use of STRIDE with high input dimension $d$ without separate feature extraction, for instance, should still be very challenging. STRIDE profits from sparse GPR, but the more flexible deep GP models still require more computational resources.
@@
S.U, A.L.T., and J.L. acknowledge the use of the HWU high-performance computing facility (DMOG) and associated support services in the completion of this work. A.L.T. was partially supported by EPSRC grants EP/X01259X/1 and EP/Y028783/1.